OpenStudy (anonymous):

is -7+9=-9+7 true, false or open?

OpenStudy (akashdeepdeb):

-7 + 9 is 2 And -9 + 7 is -2 What do you think? :)

OpenStudy (anonymous):

-7+9=2 -9+7=-2 hence false

OpenStudy (anonymous):

It false right

OpenStudy (akashdeepdeb):

Yes! You are right!! It is false! :D